# BranchGRPO: Stable and Efficient GRPO with Structured Branching in Diffusion Models

### v3

@@ Abstract
Recent progress in aligning image and video generative models with Group Relative Policy Optimization (GRPO) has improved human preference alignment, yet existing approaches still suffer from high computational cost due to sequential rollouts and large numbers of SDE sampling steps, as well as training instability caused by sparse rewards. In this paper, we present BranchGRPO, a method that restructures the rollout process into a branching tree, where shared prefixes amortize computation and pruning removes low-value paths and redundant depths. BranchGRPO introduces three contributions: (1) a branch sampling scheme that reduces rollout cost by reusing common segments; (2) a tree-based advantage estimator that converts sparse terminal rewards into dense, step-level signals; and (3) pruning strategies that accelerate convergence while preserving exploration. On HPDv2.1 image alignment, BranchGRPO improves alignment scores by up to 16% over strong baselines, while reducing per-iteration training time by nearly 55%. On WanX-1.3B video generation, it further achieves higher Video-Align scores with sharper and temporally consistent frames compared to DanceGRPO. Codes and models are available at BranchGRPO.
@@ 1 Introduction
Within RLHF, Group Relative Policy Optimization (GRPO) has shown strong stability and scalability across text-to-image and text-to-video tasks (Liu et al., 2025a; Xue et al., 2025). However, when applied to diffusion and flow-matching models, current GRPO variants still face two fundamental bottlenecks: (1) Inefficiency. Standard GRPO adopts a sequential rollout design, where each trajectory must be independently sampled under both the old and new policies. This incurs $O(N\cdot T)$ complexity with denoising steps $T$ and group size $N$, leading to significant computational redundancy and limiting scalability in large-scale image and video generation tasks. (2) Sparse rewards. Most methods compute only a single reward from the final denoised image or video and propagate it uniformly across all denoising steps. This trajectory-level supervision neglects informative signals contained in intermediate states, resulting in unstable optimization, inaccurate credit assignment, and high-variance gradients. This raises a central challenge: how can we accurately attribute sparse outcome rewards to the specific denoising steps that truly contribute to high-quality generations?
@@
Crucially, the tree structure enables a novel reward fusion and depth-wise advantage estimation mechanism. Instead of assigning a single terminal reward to every denoising step, BranchGRPO aggregates rewards from leaf nodes predicted by the reward model and propagates them backward through depth-wise normalization. This produces layer-wise advantages that allow for finer-grained credit assignment and more stable policy optimization. In addition, BranchGRPO integrates width- and depth-level pruning strategies that remove low-value or redundant branches. These pruning mechanisms accelerate training, reduce computational overhead, and enhance effective exploration by reallocating resources to promising regions of the trajectory space.
@@
  • We validate the effectiveness of BranchGRPO on both text-to-image and image-to-video alignment tasks, demonstrating its general applicability across modalities.
@@ 2 Related Work
More recently, Group Relative Policy Optimization (GRPO) (Shao et al., 2024) has been introduced as a scalable alternative to PPO for preference optimization. DanceGRPO (Xue et al., 2025) pioneers the application of GRPO to visual generation, unifying diffusion and flow models via SDE reformulation and demonstrating stable optimization across text-to-image, text-to-video, and image-to-video tasks. Flow-GRPO (Liu et al., 2025a) extends GRPO to flow matching models by introducing ODE-to-SDE conversion and denoising reduction for improved efficiency. MixGRPO (Li et al., 2025a) further enhances efficiency with a mixed ODE–SDE sliding-window scheme, though it still faces trade-offs between overhead and performance. Our work continues this line by introducing BranchGRPO, which leverages branching rollouts, depth-wise reward fusion, and structured pruning to improve both stability and efficiency; while related to TreePO in LLMs (Li et al., 2025b), our method adapts tree-structured rollouts specifically to diffusion dynamics.
@@ 3.1 Does Branch Rollout Harm Diversity?
[…]
@@ 3.2 Branch Rollout Algorithm
Unlike prior GRPO variants such as DanceGRPO and FlowGRPO that rely on sequential rollouts, where each trajectory is sampled independently from start to finish, BranchGRPO reorganizes the process into a tree-structured rollout (Figure 1). For each prompt, we initialize a root node with a same initial noise $z_{0}\sim\mathcal{N}(0,I)$ and then denoise step by step along the reverse SDE. At designated split steps $\mathcal{B}$, the current state expands into $K$ children, producing multiple sub-trajectories that share early prefixes but diverge afterward. The branching is achieved by injecting stochastic perturbations into the SDE transition, with a hyperparameter $s$ controlling the correlation strength among child noises. This mechanism balances exploration diversity and stability while keeping the marginal distribution unchanged. The rollout continues until reaching the maximum depth $T$, at which point all leaves are collected for reward evaluation. Formally, the baseline reverse SDE is
$\mathrm{d}z_{t}=\Big{(}f(z_{t},t)-\tfrac{1}{2}g^{2}(t)\nabla\log p_{t}(z_{t})\Big{)}\mathrm{d}t\;+\;g(t)\,\mathrm{d}w_{t},$ (1)
while our branching modification at split step $i\in\mathcal{B}$ produces
$z_{i+1}^{(b)}=\mu_{\theta}(z_{i},t_{i})+s_{i}\sqrt{h_{i}}\,\xi_{b},\qquad\xi_{b}=\frac{\xi_{0}+s\,\delta\xi_{b}}{\sqrt{1+s^{2}}},\quad b=1,\dots,K,$ (2)
where $\xi_{0},\delta\xi_{b}\sim\mathcal{N}(0,I)$ and $s\!\geq\!0$ adjusts inter-branch correlation. When $s=0$, all branches collapse to a single path, while as $s\!\to\!\infty$, the branches become independent.Intuitively, smaller $s$ yields highly correlated children (stable but limited exploration), whereas larger $s$ produces more independent trajectories (diverse but noisier).
@@ 3.3 Reward Fusion and Depth-wise Advantage Estimation
The branch rollout constructs a full trajectory tree in which all leaf nodes share common ancestral prefixes. This shared structure makes it natural to represent the rewards of intermediate nodes through their descendant leaves, analogous to how genetic traits of ancestors can be inferred from their offspring.
In contrast, existing GRPO variants update every denoising step using only the final reward, which ignores informative intermediate states and leads to unstable credit assignment with high variance (see Fig. 3). BranchGRPO leverages its tree structure to overcome this limitation: leaf rewards are propagated upward to internal nodes, normalized layer by layer, and transformed into dense step-wise advantages—without introducing any additional models or computational overhead. Our optimization objective remains a clipped GRPO loss defined on tree edges:
@@
where $\mathcal{E}$ denotes tree edges and $\rho_{e}(\theta)$ is the importance ratio.
@@
Directly averaging leaf rewards to estimate internal node values can be dominated by rare but noisy trajectories, leading to unstable gradients. To mitigate this, we fuse descendant rewards using path-probability weighting: let $r_{\ell}$ be the reward of a leaf $\ell$. For an internal node $n$ with children $\mathcal{L}(n)$, we fuse its rewards by
$\bar{r}(n)=\sum_{\ell\in\mathcal{L}(n)}w_{\ell}^{(n)}\,r_{\ell},\qquad w_{\ell}^{(n)}=\frac{\exp(\beta\,s_{\ell})}{\sum_{j\in\mathcal{L}(n)}\exp(\beta\,s_{j})},$ (4)
where $s_{\ell}$ is the path log-probability under the behavior policy and $\beta$ controls the strength of weighting. When $\beta=0$, this reduces to simple averaging; larger $\beta$ emphasizes high-likelihood paths and reduces the influence of noisy, low-probability leaves, in all our experiments we set $\beta=1.0$.
We consider two alternatives for aggregating leaf rewards. Uniform averaging treats all leaves equally is robust to log-prob calibration errors and encourages exploration by preserving contributions from low-probability leaves. However, it deviates from the expected return under the behavior policy and can introduce high variance when the number of branches is large.
In contrast, path-weighted scheme aligns node-level credit assignment with the behavior distribution, typically reducing variance and yielding more stable updates. Yet, it relies on log-prob calibration and may overly concentrate weight on high-likelihood paths in deep trees. We empirically compare both variants (Sec. 4.4), showing that path weighting improves stability in practice, while uniform averaging provides stronger exploratory diversity.
@@
The tree depth corresponds to denoising steps, and nodes at the same depth share the same noise level. This makes rewards across nodes within a depth directly comparable, while rewards across different depths may differ drastically in scale due to varying noise states. To ensure fair comparison, we normalize aggregated rewards separately within each depth $d$:
@@
where $\mathcal{N}_{d}$ denotes the set of nodes at depth $d$. Each edge advantage $A(e)$ is then inherited from its child node and optionally clipped to $A_{\max},A_{\max}]$. This per-step standardization aligns gradient magnitudes with the denoising process and prevents deeper nodes from being dominated by shallow ones.
Compared to GRPO’s prompt-level normalization, which uniformly broadcasts a single terminal reward, our branch-based scheme yields process-dense credit signals that respect the denoising trajectory. By normalizing rewards at the step level, BranchGRPO produces more stable gradients and more accurate credit assignment to the intermediate denoising steps that truly matter.
@@ 3.4 Pruning Strategies
While branch rollouts improve efficiency and provide dense process rewards, an excessive number of branches may induce exponential growth in trajectory count, leading to prohibitive backpropagation cost. Recent work (Li et al., 2025a) has revealed that standard GRPO training contains substantial redundancy and pruning potential. Motivated by this, we introduce two complementary pruning strategies in the context of BranchGRPO: width pruning, which reduces the number of leaves used for gradient updates, and depth pruning, which skips unnecessary denoising layers.
Importantly, pruning is applied only after reward fusion and depth-wise normalization, and affects backpropagation but not forward rollouts or reward evaluation. This design ensures that all trajectories contribute to reward signals, while gradients are computed only for the selected subset of nodes.
@@
After computing rewards and advantages for all leaves $\mathcal{L}$, we optionally select a subset of them for gradient updates. We study two simple modes. The first is Parent-Top1, which retains the higher-reward child of each parent in the last branching step. This strategy halves gradient cost while ensuring coverage of all parents, yielding stable but slightly less diverse updates. The second is Extreme selection, which preserves both the globally best and worst $b$ leaves by reward score. By explicitly maintaining strong positive and negative signals, this mode injects more exploratory feedback but introduces higher variance.
@@
Notably, recent work MixGRPO (Li et al., 2025a) achieves strong performance by hybrid ODE–SDE, effectively acting as a form of depth pruning. This motivates our explicit exploration of depth pruning within BranchGRPO.
Specifically, we maintain a set of pruned depths $\mathcal{D}$ and ignore gradients from nodes at these layers while still performing full forward rollouts for reward evaluation. To avoid permanently excluding certain steps, we employ a sliding window: the pruned depths shift gradually toward deeper layers as training progresses, until reaching a predefined maximum depth. Formally, pruning is enabled in all of training steps, and every fixed interval the pruning window slides one step deeper, until reaching a pre-defined stop depth.
This dynamic scheme improves efficiency while still accounting for the contributions of gradients from different depths, thus balancing computational savings with policy effectiveness.
@@ 4.2 Implementation Details
We set the tree depth to $d=20$ and the branch factor to $b=2$, yielding $16$ leaves per rollout before pruning. The branching steps $\mathcal{B}$ use three presets: Dense $(0,3,6,9)$ as the default, Mixed $(0,5,15,20)$, and Sparse $(0,10,13,19)$. The branch correlation sweeps $s\in\{0,1,2,4,8\}$. Training runs for 300 optimizer steps with gradient accumulation $g=12$ and per-GPU batch size =2=2, on $16\times$ NVIDIA H20z GPUs. Optimization uses AdamW (learning rate $1\times 10^{-5}$, weight decay $1\times 10^{-4}$) with bf16 precision and EMA weights stored on CPU. All GRPO-related hyperparameters are kept identical across methods, with full details deferred to the supplementary material.
@@
Table 1: Efficiency–quality comparison under identical data, rewards, and optimization schedules. Best and second-best per column are in bold and underline. Speedup is relative to DanceGRPO (tf=1.0). NFE counts the expected number of denoising steps where the policy is updated; for branching methods this is an effective NFE that reflects split density and shared prefixes.
@@ 4.3 Main Results
Table 1 compares efficiency and alignment quality across methods, while Figure 1 shows reward dynamics during training. BranchGRPO achieves consistently better alignment than DanceGRPO while using substantially fewer function evaluations and shorter iteration times. MixGRPO improves efficiency but at the cost of weaker alignment, whereas BranchGRPO provides a more favorable balance, and pruning variants further reduce computation without hurting performance.
The reward curves in Figure 1 also highlight the stability of our approach. DanceGRPO with reduced timestep fractions suffers from strong reward fluctuations, while the full-timestep variant converges more smoothly but at much higher cost. BranchGRPO combines the best of both, offering faster reward growth and more stable trajectories. Figure 4 further illustrates qualitative improvements: compared with Flux and DanceGRPO, our method produces more faithful, detailed, and coherent generations across diverse prompts.
Taken together, these results demonstrate that BranchGRPO is both more efficient and more stable than existing GRPO variants, yielding superior preference alignment under significantly reduced computational overhead.

### v4

@@ Abstract
Recent progress in aligning image and video generative models with Group Relative Policy Optimization (GRPO) has improved human preference alignment, but existing variants remain inefficient due to sequential rollouts and large numbers of sampling steps, unreliable credit assignment: sparse terminal rewards are uniformly propagated across timesteps, failing to capture the varying criticality of decisions during denoising. In this paper, we present BranchGRPO, a method that restructures the rollout process into a branching tree, where shared prefixes amortize computation and pruning removes low-value paths and redundant depths. BranchGRPO introduces three contributions: (1) a branching scheme that amortizes rollout cost through shared prefixes while preserving exploration diversity; (2) a reward fusion and depth-wise advantage estimator that transforms sparse terminal rewards into dense step-level signals; and (3) pruning strategies that cut gradient computation but leave forward rollouts and exploration unaffected. On HPDv2.1 image alignment, BranchGRPO improves alignment scores by up to 16% over DanceGRPO, while reducing per-iteration training time by nearly 55%. A hybrid variant, BranchGRPO-Mix, further accelerates training to 4.7× faster than DanceGRPO without degrading alignment. On WanX video generation, it further achieves higher Video-Align scores with sharper and temporally consistent frames compared to DanceGRPO. Codes are available at BranchGRPO.
@@ 1 Introduction
Within RLHF, Group Relative Policy Optimization (GRPO) has shown strong stability and scalability across text-to-image and text-to-video tasks (Liu et al., 2025a; Xue et al., 2025). However, when applied to diffusion and flow-matching models, current GRPO variants still face two fundamental bottlenecks: (1) Inefficiency. Standard GRPO adopts a sequential rollout design, where each trajectory must be independently sampled under both the old and new policies. This incurs $O(N\cdot T)$ complexity with denoising steps $T$ and group size $N$, leading to significant computational redundancy and limiting scalability in large-scale image and video generation tasks. (2) Sparse rewards. Existing methods assign a single terminal reward uniformly across all steps, neglecting informative signals from intermediate states. This uniform propagation leads to unreliable credit assignment and high-variance gradients, raising the central question: how can we attribute sparse outcome rewards to the specific denoising steps that truly shape final quality?
@@
The tree structure further enables a novel reward fusion with depth-wise advantage estimation. Instead of uniformly propagating a single terminal reward, BranchGRPO aggregates leaf rewards and propagates them backward with depth-wise normalization, producing finer-grained step-level advantages. In addition, width- and depth-pruning strategies remove redundant branches during backpropagation, accelerating training and reallocating computation toward promising regions of the trajectory space.
@@
We validate BranchGRPO on both text-to-image and image-to-video alignment tasks, demonstrating its effectiveness and generality across modalities.
@@ 2 Related Work
More recently, Group Relative Policy Optimization (GRPO) (Shao et al., 2024) has been introduced as a scalable alternative to PPO for preference optimization. DanceGRPO (Xue et al., 2025)and Flow-GRPO (Liu et al., 2025a) pioneers the application of GRPO to visual generation, unifying diffusion and flow models via SDE reformulation and demonstrating stable optimization across text-to-image, text-to-video, and image-to-video tasks. TempFlow-GRPO (He et al., 2025) further highlights the limitation of sparse terminal rewards with uniform credit assignment, proposing temporally-aware weighting across denoising steps. MixGRPO (Li et al., 2025a) further enhances efficiency with a mixed ODE–SDE sliding-window scheme, though it still faces trade-offs between overhead and performance. Our work continues this line by introducing BranchGRPO, which leverages branching rollouts, depth-wise reward fusion, and structured pruning to improve both stability and efficiency; while related to TreePO in LLMs (Li et al., 2025b), our method adapts tree-structured rollouts specifically to diffusion dynamics.
@@ 3.1 Does Branch Rollout Harm Diversity?
Figure 2: Comparison of sequential and branch rollouts. Left/Right: example generations from DanceGRPO and BranchGRPO, respectively. Middle: distribution of sampled images projected into 2D feature space, where red and blue dots correspond to DanceGRPO and BranchGRPO.
A natural concern with branch rollouts is that reusing shared prefixes might reduce sample diversity. To investigate this, we generate 4096 samples each from DanceGRPO and BranchGRPO and evaluate their distributions across multiple feature spaces. Figure 2 provides both qualitative and quantitative evidence: the left/right panels show representative generations, while the middle panel visualizes the sampled distributions in a 2D feature embedding, where DanceGRPO and BranchGRPO points largely overlap.
@@
Taken together, Figure 2 and these statistics demonstrate that branch rollouts preserve distributional and semantic diversity, introducing at most negligible shifts across different feature spaces.
@@ 3.2 Branch Rollout Algorithm
Unlike prior GRPO variants such as DanceGRPO and FlowGRPO that rely on sequential rollouts, where each trajectory is sampled independently from start to finish, BranchGRPO reorganizes the process into a tree-structured rollout (Figure 1). For each prompt, we initialize a root node with a same initial noise $z_{0}\sim\mathcal{N}(0,I)$ and then denoise step by step along the reverse SDE. At designated split steps $\mathcal{B}$, the current state expands into $K$ children, producing multiple sub-trajectories that share early prefixes but diverge afterward. The branching is achieved by injecting stochastic perturbations into the SDE transition, with a hyperparameter $s$ controlling the correlation strength among child noises. This mechanism balances exploration diversity and stability while keeping the marginal distribution unchanged. The rollout continues until reaching the maximum depth $T$, at which point all leaves are collected for reward evaluation.
Formally, following Xue et al. (2025), the reverse-time dynamics can be written in the unified SDE form:
$\mathrm{d}z_{t}\;=\;\Big{(}f_{t}z_{t}-\tfrac{1+\varepsilon_{t}^{2}}{2}\,g_{t}^{2}\nabla\log p_{t}(z_{t})\Big{)}\,\mathrm{d}t\;+\;\varepsilon_{t}g_{t}\,\mathrm{d}w_{t},$ (1)
where $\varepsilon_{t}$ controls stochasticity.
At a split step $i\in\mathcal{B}$ with step size $h_{i}=t_{i}-t_{i+1}$, instead of sampling a single successor we generate $K$ correlated children:
$z_{i+1}^{(b)}=\mu_{\theta}(z_{i},t_{i})\;+\;\varepsilon_{t_{i}}g_{t_{i}}\sqrt{h_{i}}\,\xi_{b},\qquad\xi_{b}=\frac{\xi_{0}+s\,\eta_{b}}{\sqrt{1+s^{2}}},\quad b=1,\dots,K,$ (2)
where $\xi_{0},\eta_{b}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,I)$, with $\xi_{0}$ shared across branches and $\eta_{b}$ denoting branch-specific innovations. The parameter $s\!\geq\!0$ tunes inter-branch correlation: small $s$ yields highly correlated but stable branches, while large $s$ makes branches nearly independent. By construction, each $\xi_{b}\sim\mathcal{N}(0,I)$, so every child $z_{i+1}^{(b)}$ has the same marginal distribution as the baseline SDE step.
@@ 3.3 Reward Fusion and Depth-wise Advantage Estimation
Branch rollouts construct a trajectory tree where all leaves share ancestral prefixes. This shared structure enables internal node values to be represented by their descendant rewards, naturally allowing backward propagation of leaf signals.
In contrast, existing GRPO variants update every step using only a single terminal reward, ignoring intermediate states and resulting in unreliable credit assignment with high variance (Fig. 3). BranchGRPO mitigates this limitation by propagating leaf rewards upward, fusing them through path-probability weighting, and applying depth-wise normalization to produce dense step-level advantages. The optimization objective remains the clipped GRPO loss defined on tree edges:
@@
where $\mathcal{E}$ denotes tree edges and $\rho_{e}(\theta)$ the importance ratio.
@@
Directly averaging leaf rewards may be dominated by rare noisy trajectories. We instead apply a weighted fusion: for internal node $n$ with descendant leaves $\mathcal{L}(n)$,
$\bar{r}(n)=\sum_{\ell\in\mathcal{L}(n)}w_{\ell}^{(n)}\,r_{\ell},\qquad w_{\ell}^{(n)}=\frac{\exp(\beta s_{\ell})}{\sum_{j\in\mathcal{L}(n)}\exp(\beta s_{j})},\;\;s_{\ell}=\log p_{\text{beh}}(\ell\mid n).$ (4)
Here $p_{\text{beh}}$ is the behavior policy and $\beta$ controls concentration. $\beta=0$ reduces to uniform averaging; $\beta=1$ recovers self-normalized importance weighting aligned with the behavior distribution. Uniform averaging is robust to log-prob calibration errors and encourages exploration by retaining low-probability leaves, but introduces variance when branches are many. Path-weighting reduces variance and stabilizes updates, though it may over-concentrate on high-likelihood leaves in deep trees. We empirically compare both variants in Sec. 4.4.
@@
Nodes at the same depth share the same noise level and are thus directly comparable, while rewards across depths vary drastically due to changing noise states. To balance gradient contributions, we normalize aggregated rewards within each depth $d$:
@@
where $\mathcal{N}_{d}$ denotes all nodes at depth $d$. Each edge advantage $A(e)$ inherits from its child node and is optionally clipped to $A_{\max},A_{\max}]$. This per-depth standardization prevents late denoising steps with smaller variance from dominating, yielding process-dense and balanced credit signals. Compared to GRPO’s prompt-level normalization, which broadcasts a single terminal reward, our scheme produces stable gradients and finer credit assignment to the denoising steps that matter.
@@ 3.4 Pruning Strategies
While branch rollouts improve efficiency and provide dense process rewards, an excessive number of branches may induce exponential growth in trajectory count, leading to prohibitive backpropagation cost. To further accelerate training, we introduce two complementary pruning strategies in the context of BranchGRPO: width pruning, which reduces the number of leaves used for gradient updates, and depth pruning, which skips unnecessary denoising steps.
Importantly, pruning is applied only after reward fusion and depth-wise normalization, and affects backpropagation but not forward rollouts or reward evaluation. This design ensures that all trajectories contribute to reward signals, while gradients are computed only for the selected subset.
@@
After computing rewards and advantages for all leaves $\mathcal{L}$, we restrict gradient updates to a subset of them. We investigate two modes. The first, Parent-Top1, retains the child with the higher reward from each parent at the last branching step. This strategy roughly halves gradient computation while ensuring coverage of all parents, yielding stable but slightly less diverse updates. The second, Extreme selection, preserves both the globally best and worst $b$ leaves by reward score. This explicitly maintains strong positive and negative signals, which may enhance exploration but also increase variance.
@@
Branch rollouts generate dense rewards across all denoising steps, but computing gradients at every depth remains costly. To improve efficiency, we introduce depth pruning, which skips gradient computation at selected timesteps . Concretely, we maintain a set of pruned depths $\mathcal{D}$ and ignore gradients from nodes at these layers. To prevent permanently discarding certain steps, we adopt a sliding window schedule: the pruned depths gradually shift toward later timesteps as training progresses, until reaching a predefined maximum depth. Formally, pruning is active throughout training, and at fixed intervals the window slides one step deeper until the stop depth is reached.
@@ 4.2 Implementation Details
We set the tree depth to $d=20$ and the branch factor to $b=2$, yielding $16$ leaves per rollout before pruning. The branching steps $\mathcal{B}$ use three presets: Dense $(0,3,6,9)$ as the default, Mixed $(0,5,15,20)$, and Sparse $(0,10,13,19)$. The branch correlation sweeps $s\in\{0,1,2,4,8\}$. Training runs for 300 optimizer steps with gradient accumulation $g=12$ and per-GPU batch size =2=2, on $16\times$ NVIDIA H200 GPUs. Optimization uses AdamW (learning rate $1\times 10^{-5}$, weight decay $1\times 10^{-4}$) with bf16 precision and EMA weights stored on CPU. All GRPO-related hyperparameters are kept identical across methods, with full details deferred to the supplementary material.
@@
Table 1: Efficiency–quality comparison. The best and second-best results in each column are highlighted in bold and underline, respectively. NFE denotes the number of function evaluations of the denoiser. For branching methods, we report the average per-sample NFE, computed as the total function evaluations in the tree divided by the number of final samples.
@@
BranchGRPO-Mix 13.68 4.25 148 0.363 0.230 1.290
@@ 4.3 Main Results
Table 1 compares efficiency and alignment quality. BranchGRPO achieves superior alignment while substantially reducing cost: for example, BranchGRPO-DepPruning improves HPS-v2.1 from 0.360 (DanceGRPO) to 0.369 (+2.5%) and attains the best PickScore (0.231) and ImageReward (1.319). Meanwhile, BranchGRPO-Mix accelerates training to 148s per iteration—over $4.7\times$ faster than DanceGRPO(tf=1.0)—while maintaining competitive alignment scores. MixGRPO also reduces computation (289s) but its alignment scores plateau at DanceGRPO levels.
The reward curves in Figure 1 further highlight stability. DanceGRPO(tf=0.6) suffers from high-variance fluctuations, while the full-timestep variant converges more smoothly but at much higher cost. BranchGRPO combines the advantages of both, yielding faster early reward growth and smoother convergence. Qualitative comparisons in Figure 4 show that our generations exhibit sharper details and stronger semantic alignment compared with Flux and DanceGRPO.